A 2.00 g sample of an iron ore is dissolved in acid and then reduced to Fe²⁺. It was then titrated against 0.025 M KMnO₄ solution and used 47.5 mL to reach the end point. Calculate the %Fe₃O₄ (231.54 g/mol) of the ore sample.



1
Expert's answer
2022-01-18T03:25:05-0500

Solution:

Balanced chemical equation: Fe3O4 → 3Fe2+

1 Fe3O4 ≡ 3 Fe2+

Balanced chemical equation: 5Fe2+ + MnO4 + 8H+ → 5Fe3+ + Mn2+ + 4H2O

5 Fe2+ ≡ 1 MnO4

Therefore,

5 Fe3O4 ≡ 15 Fe2+ ≡ 3 MnO4

and

stoichiometric ratio = 5 mol Fe3O4 / 3 mol MnO4


Thus,

(0.0475 L KMnO4) × (0.025 mol KMnO4 / 1 L KMnO4) × (5 mol Fe3O4 / 3 mol KMnO4) = 0.002 mol Fe3O4

The molar mass of Fe3O4 is 231.54 g/mol

Therefore,

(0.002 mol Fe3O4) × (231.54 g Fe3O4 / 1 mol Fe3O4) = 0.463g Fe3O4


%Fe3O4 = (Mass of Fe3O4 / Mass of sample) × 100%

%Fe3O4 = (0.463 g / 2.00 g) × 100% = 23.15%

%Fe3O4 = 23.15%


Answer: %Fe3O4 = 23.15%